OpenStudy (anonymous):

If u were to use the substitution method to solve the following system, choose the new equation after the expression equivalent to x from the 2nd equation is substituted into the 1st equation. 2x – 3y = –29 x + 4y = 13

OpenStudy (anonymous):

x+4y=13 x=13-4y substitute x=13-4y into 2x-3y=-29 2(13-4y)-3y=-29 26-8y-3y=-29 -11y=-29-26 y=55/11 y=5 using y=5, substitute it into x=13-4y x=13-4(5) x=13-20 x=-7

OpenStudy (anonymous):

Do you mean it'd be 2x – 3(–4y + 13) = –29??

OpenStudy (anonymous):

@monokerous?!

OpenStudy (anonymous):

no. it's correct. x=13-4y so if you plug this into the equation, 2x becomes 2(13-4y)